OpenStudy (anonymous):

Which of the following represents fifth root of x squared in exponential form? x to the 2 fifths power x to the 5 over 2 power 5x2 2x5

OpenStudy (anonymous):

the exponent should be \[\large \frac{2}{5}\]

jimthompson5910 (jim_thompson5910):

Using the rule \[\Large \sqrt[n]{x^m} = x^{m/n}\] we can say \[\Large \sqrt[5]{x^2} = x^{2/5}\]